# The Structure of Financial Equity Research Reports – Identification of the Most Frequently Asked Questions in Financial Analyst Reports to Automate Equity Research Using Llama 3 and GPT-4

### v1

@@ 4.3 Qualitative Validation of the Question List Using Expert Interviews and Prior Research
Figure 6: Correctness of Answers by GPT-4-turbo-2024-04-09, Llama-3-70B, and the Best of Both Models. The green parts show correct answers, the red-white-hatched parts show errors. Interestingly, the errors of GPT-4 and Llama 3 have almost no overlap. When one model is unable to correctly answer a question, the other model usually is. There was only one question that both models did not answer correctly despite the relevant information being present in the prompt. The fine black lines in the right plot delimit different questions, and the culmination of errors for certain questions shows that the models have difficulty with particular questions.
As a qualitative side note, we found that GPT-4 tends to provide longer responses with more context. In some cases, we thus found that GPT-4 provided helpful context that Llama-3-70B missed. We performed some follow-up tests with Llama 3 to see if this difference was a lack of capability or simply a difference in the default verbosity among the models. We found that, for the purpose of financial text, Llama 3 is able to provide the same extraction depth and abstraction capabilities as GPT-4 does. But Llama 3 tends to provide more direct answers compared to GPT-4 unless prompted to add contextual flavor. Also, Llama often attempts to calculate growth numbers when asked about rates. In all attempts, it fails to provide correct absolute or relative year-on-year changes, but stays in the correct ballpark of plus/minus 10%.
The language models correctly identified information in full-form text and in tabular format. Tables were simply copy-pasted from the annual reports, so that the formatting of these tables was not specifically optimized language model readability. The models showed high robustness in extracting financial information from tables. For each mistake made by the language models, we inspected the context to see if a human had been able to answer the question given the text-only context (no PDF formatting was provided, limiting what the language models were able to see relative to what a human would be able to infer from a text’s context in the annual report). We made sure that no language model answer was marked as incorrect if there was no clear answer in the context, but no such cases occurred in the sample of 200 questions.
5 Results
@@
5.1 Result Overview
In summary, 78.7% of the 169 questions in ERRs are automatable. 48.2% of the statements in ERRs are extractable. 30.4% of questions require access to non-public databases but have potential for automation. Only 21.3% of questions require judgment that goes beyond extraction.
5.2 Analysis of Question Categories and Subcategories
Table 1: Question Frequency List of the Five Most Frequently Answered Questions. The table indicates which types of statements are the most frequent across all ERRs. Across the 72 reports, 66 contained information about the stock price and 65 about the company’s challenges. Company details and market environment statements appear less frequently.
# Question Count Subcategory Numerical
@@ 5.2 Analysis of Question Categories and Subcategories
1 Stock price 66 Stock - Price Yes No
2 Potential challenges 65 Company - Other No No
3 Primary industry 64 Product No Yes
4 Analyst score 64 Financials - Other No Yes
5 Earnings per share 62 Financials - Financing Yes Yes
A share of 73.4% of statements in the category Product are automatable. The most critical question category by statement count is the Financials. 70.6% of statements from this category are extractable. A share of 54.6% is automatable in the category Company. A share of 16.6% is automatable in the category Stock. A share of 4.6% is automatable in the category Market. This is because statements about the market environment usually require access to diverse sources outside the company’s annual and quarterly reports. None of the statements from the Analysis category can be automated with extractions from publicly available corporate reports. This category contains the target price (forward guidance), recommendation, and risk assessments.
5.3 Analysis of Extractable and Non-Extractable Statements
Table 2: Overview of Extractable and Non-extractable Statements. Numeric-extractable statements have a share of 40%. Numeric-non-extractable statements have a share of 26%. Almost two-thirds of non-numeric statements are non-extractable.
@@ 5.3 Analysis of Extractable and Non-Extractable Statements
Table 2 portrays the classifications of questions answered in the examined ERRs. The x-axis shows whether the questions answered are numeric or non-numeric. The y-axis exhibits whether the information gained is extractable from another source of non-written text.
@@
Two-thirds are numeric, and more than half are extractable. While extractable information is mostly numeric (40%), extractable non-numeric information is rare (12%). Out of the non-extractable information, slightly more is numeric.
5.4 Analysis of Contextualizing and Summarizing Components
The Analysis question category contains mostly summarizing and contextualizing components. Given the same set of facts, different analysts may weigh, select, and combine those facts differently, leading to different recommendations.
Related to this, there are numerous questions in the category Market of this nature. Market developments require simplification and curation to distill into a few pages of text. Similar to the Analysis category, different observers judge the same set of facts differently, leading to different conclusions. Given that the potential inputs to this category are vast, with many news reports and other sources to choose from, it is unlikely that an automated system can already handle this task.
5.5 Tabular Data
Most statements in ERRs are textual or tabular. Chen et al., (2021) show that numerical reasoning across tables and text is feasible (pp. 5–7). Zhu et al., (2021) confirm that tabular information extraction is possible, particularly for financial data (pp. 3282–3284).
We confirm these findings: Our validation from section 4.4 has not required any manual formatting of table data – we copied tables from annual reports without formatting into the models’ context, and they extracted information from these ill-formatted strings with high reliability.
6 Conclusion
Our results confirm the findings by Coleman et al., (2022) that partly automating equity research reports (ERRs) is feasible. Only 21.3% of questions require complex judgment that takes into consideration more information than would fit in a language models’ context window.
In particular, the Analysis category in ERRs contains summarizing statements that make recommendations. These statements are not extractable from other sources. Only 3.5% of statements in ERRs fall under the analysis category. Given the oversized importance of this category, and given that humans may, for the time being, still be better at providing high-stake recommendations, this category should be left to human financial analysts.
@@ 6 Conclusion
Figure 2 shows that, without weighting the questions by their occurrence frequency, 48.2% of questions answered in ERRs can be answered by extracting information from public textual sources. An additional 30.4% are available in proprietary market data sources.
@@
Direct information extraction from ERRs is still a largely unexplored field. Various approaches were presented in section 2.1. Future research can implement these methods.

### v2

@@ IV-C Qualitative Validation of the Question List Using Expert Interviews and Prior Research
Figure 6: Correctness of Answers by GPT-4-turbo-2024-04-09, Llama-3-70B, and the Best of Both Models. The green parts show correct answers, the red-white-hatched parts show errors. Interestingly, the errors of GPT-4 and Llama 3 have almost no overlap. When one model is unable to correctly answer a question, the other model usually is. There was only one question that both models did not answer correctly despite the relevant information being present in the prompt. The fine black lines in the second plot delimit different questions, and the culmination of errors for certain questions shows that the models have difficulty with particular questions.
As a qualitative side note, we found that GPT-4 tends to provide longer responses with more context. In some cases, we thus found that GPT-4 provided helpful context that Llama-3-70B missed. We performed some follow-up tests with Llama 3 to see if this difference was a lack of capability or simply a difference in the default verbosity among the models. We found that, for the purpose of financial text, Llama 3 is able to provide the same extraction depth and abstraction capabilities as GPT-4 does. But Llama 3 tends to provide more direct answers compared to GPT-4 unless prompted to add contextual flavor. Also, Llama often attempts to calculate growth numbers when asked about rates. In all attempts, it fails to provide correct absolute or relative year-on-year changes but stays in the correct ballpark of plus/minus 10%.
Llama 3 was trained on sequences of 8,192 tokens. Annual reports are usually much longer, often having around 100,000 tokens.
The language models correctly identified information in full-form text and in tabular format. Tables were simply copy-pasted from the annual reports, so the formatting of these tables was not specifically optimized for language model readability. The models showed high robustness in extracting financial information from tables. For each mistake made by the language models, we inspected the context to see if a human had been able to answer the question given the text-only context (no PDF formatting was provided, limiting what the language models were able to parse relative to what a human would be able to visually infer from the format in the annual report). We made sure that no language model answer was marked as incorrect if there was no clear answer in the context, but no such cases occurred in the sample of 200 questions.
V Results
@@
V-A Result Overview
In summary, 75.15% of the 169 questions in ERRs are automatable. More precisely, 51.91% of the statements in ERRs are extractable, and 24.24% of questions require access to non-public databases but have potential for automation. Only 24.85% of questions require judgment that goes beyond extraction from either a corporate report or from a financial database.
V-B Analysis of Question Categories and Subcategories
TABLE I: Question Frequency List of the Five Most Frequently Answered Questions. The table indicates which types of statements are the most frequent across all ERRs. Across the 72 reports, 66 contained information about the stock price and 65 about challenges. Company details and market environment statements appear less frequently.
Question Count Subcategory Numerical
@@ V-B Analysis of Question Categories and Subcategories
Key financials 122 Financials - Other Yes Yes
Analyst rating 64 Analysis - Recommend. No No
Cash flow 64 Financials - P&L Yes Yes
Target price? 62 Analysis - Recommend. Yes No
Revenue over time 60 Financials - P&L Yes Yes
A share of 73.4% of statements in the category Product are automatable. Financials is the most critical question category by statement count. 70.6% of statements from this category are extractable. A share of 54.6% is automatable in the category Company. A share of 16.6% is automatable in the category Stock. A share of 4.6% is automatable in the category Market. This is because statements about the market environment usually require access to diverse sources outside the company’s annual and quarterly reports. None of the statements from the Analysis category can be automated with extractions from publicly available corporate reports. This category contains the target price (forward guidance), recommendation, and risk assessment.
V-C Analysis of Extractable and Non-Extractable Statements
TABLE II: Overview of Extractable and Non-extractable Statements. The table portrays the classifications of questions answered in the examined ERRs. The columns show whether the questions answered are numeric or non-numeric. The rows indicate whether the information is extractable from textual sources (such as annual reports).
@@ V-C Analysis of Extractable and Non-Extractable Statements
The first part of table II contains the classification of the 165 unique questions from all ERRs. Two-thirds of the questions are numeric, and more than half are extractable. While extractable information is mostly numeric (40.61% of total questions are numeric-extractable, 61.47% of numeric questions are extractable), extractable non-numeric information is rare (10.3% of total questions, 30.36% of non-numeric questions are extractable). Out of the non-extractable information, slightly more is numeric, but the number of numeric questions is higher (109 unique questions) than the number of non-numeric questions (56 unique questions).
The Analysis category requires special mention as it contains summarizing statements that make recommendations. These statements are not extractable from anywhere, as they require comprehension across multiple sources. Only 3.64% of statements (non-unique) in ERRs fall under the analysis category.
@@
Table II shows that, without weighting the questions by their occurrence frequency, 50.91% (40.61% + 10.3%) of questions answered in ERRs can be answered by extracting information from public textual sources.
V-D Analysis of Contextualizing and Summarizing Components
The Analysis question category contains summarizing and contextualizing components. Given the same set of facts, different analysts may weigh, select, and combine those facts differently, leading to different recommendations.
Related to this, there are numerous questions in the category Market of similar nature. Market developments require simplification and curation to distill into a few pages of text. Similar to the Analysis category, different observers judge the same set of facts differently, leading to different conclusions. Given that the potential inputs to this category are vast, with many news reports and other sources to choose from, it is unlikely that an automated system can already handle this task.
V-E Tabular Data
Most statements in ERRs are textual or tabular. [22] show that numerical reasoning across tables and text is feasible (pp. 5–7). [36] confirm that tabular information extraction is possible, particularly for financial data (pp. 3282–3284).
We confirm these findings: Our validation from section IV-D has not required any manual formatting of table data – we copied tables from annual reports without formatting into the models’ context, and they extracted information from these ill-formatted strings with high reliability.
VI Conclusion
TABLE III: List of Research Providers, Sorted by the Number of Annual Reports Used in This Study. J.P. Morgan provided the most equity research reports for this analysis with 16 pieces, followed by Deutsche Bank (9), Zacks (8), and Barclays (7). The average number of statements is 68.6, the median is 70, and the minimum number is a result of Veritas Investment Research’s one-pager with only nine statements.
Research Provider
Research Report
Counts
Avg. No. Statements
per Report
J.P. Morgan 16 93
Deutsche Bank 9 56
Zacks 8 90
Barclays 7 63
Mizuho 5 48
Needham 5 71
KBW 3 55
Refinitiv 2 53
New Constructs 2 40
Phillip Securities Res. 2 78
GlobalData 1 115
China Renaissance 1 106
IBM Res. 1 34
Punto Casa de Bolsa 1 24
Spartan Capital 1 58
Thompson Res. 1 34
Mitsubishi UFJ M.S. 1 31
Oppenheimer 1 44
BPC Res. 1 44
Veritas Investment 1 9
Echelon 1 59
finnCap 1 75
BTIG 1 69
VI-A Summary
Our results confirm the findings by [2] that partly automating equity research reports (ERRs) is feasible. Only one quarter of questions require complex judgment that takes into consideration more information than would fit in a language model’s context window.
Given the oversized importance of the Analysis category, and given that humans may still be better at providing high-stake recommendations, this category may be hard to automate with current models and be left to human financial analysts. It constitutes 3.64% of all statements of ERRs.
Another finding is that model errors often do not overlap (figure 6). Language models for information extraction show promising performance for extracting financial data. As this data is relevant to ERRs, the partial automation of ERRs appears feasible, especially when ensemble models are used that have independent blind spots.
VI-B Limitations
@@
VI-C Future Work
Direct information extraction for ERRs is still a largely unexplored field. Various technical methods were presented in the literature section II-A. Future research can implement these methods to write ERRs automatically.
Future research can produce benchmarks for ERR generation and add those to existing language model evaluation suites, adding to prior work [37, 38, 36, 22] by including very long contexts with raw annual report text.
In addition to creating suitable benchmarks, future research can develop domain-configured models that can generate ERRs from realistic sources for financial information (such as annual reports and quarterly reports). Human evaluators or standardized benchmarks can access the performance of such models.